OpenStudy (anonymous):

Help please?? Rewrite the following equation in standard form. y = 8/9x - 10 I know it is probably super easy, but I always need to solve a couple of problems before I get the hang of it again. :)

OpenStudy (raden):

y = (8/9)x - 10 first, added by 10 to both sides giving us : y + 10 = (8/9)x - 10 + 10 y + 10 = (8/9)x then multiply by 9 to both sides : 9 * (y+10) = 9 (8/9)x 9y + 90 = 8x the finally, subtracted by 9y to both sides : 9y + 90 - 9y = 8x - 9y 90 = 8x - 9y or 8x - 9y = 90 it is the standard form of straigh line equation

OpenStudy (anonymous):

Thank you so much! You made everything so much clearer! Thanks again! :)

OpenStudy (raden):

you're welcome :)